### v3

# Entanglement, Trace Anomaly, and Confinement in QCD

Abstract
We formulate confinement in QCD as an entropic surface phenomenon. Quark and gluon quantum information is localized on a transverse, entangling two‑sphere of radius $R_{\!EE}$; at this radius the QCD vacuum—partitioned by a hadron into interior and exterior regions—reaches its maximal entanglement entropy. Lattice‑QCD determinations of the scalar (trace) gravitational form factors fix both $R_{\!EE}$ and the transverse trace‑anomaly density $\rho_{h}(R_{\!EE})$, yielding a parameter‑free slope $c_{h}=8\pi^{2}R_{\!EE}^{2}\rho_{h}(R_{\!EE})$ and a mechanical entropy $S_{\!EE}(y)=c_{h}\,y$ that grows linearly with rapidity $y$. The entropy gradient $\partial_{R}S_{\!EE}$ changes sign at $R_{\!EE}$: it pushes colored degrees of freedom outward for $r<R_{\!EE}$ and pulls them inward for $r>R_{\!EE}$, thereby localizing them on the codimension‑2 entangling two‑sphere $\Sigma_{\perp}=S^{2}_{R_{\!EE}}$ (which, in the infinite‑momentum frame (IMF), projects onto the transverse plane)—the “information wall.” This provides a high‑energy (large‑$y$) entropic confinement diagnostic that complements—rather than replaces—Wilson’s area‑law criterion, which probes long‑distance dynamics near the rest frame ($y\!\to\!0$). Imposing unitarity on an entropic ansatz for the amplitude yields $\sigma(s)\!\propto\!y^{\delta}$. World data favor $\delta=2$ for elastic $pp\,(p\bar{p})$ scattering and heavy‑quark photoproduction, whereas $\phi$ photoproduction favors a softer $\delta=0.387$. All extracted cross sections remain well below the Froissart–Martin bound. These results provide a confinement criterion quantified directly from non‑perturbative QCD inputs, unifying the trace anomaly, entanglement entropy, and high‑energy scattering within a single quantitative framework.
Introduction.— Why do colored quarks and gluons never emerge from hadrons? The textbook answer invokes an area law for large Wilson loops, which diagnoses a linear confining potential in the IR [1]. Yet that criterion carries no information about ultraviolet (UV) dynamics, where asymptotic freedom reigns.
Here we propose a complementary, UV–sensitive diagnostic: entropic confinement. Quantum information carried by quarks and gluons localizes on a codimension‑2 entangling surface of fixed radius $R_{\!EE}$. Geometrically, the entangling surface is the two–sphere $\Sigma_{\perp}=S^{2}_{R}$ equipped with the induced metric $h_{ab}$ and area element $\sqrt{h}\,d^{2}x$. The universal contribution to the von Neumann entropy of the reduced state is then controlled by the QCD trace anomaly on $\Sigma_{\perp}$ [2, 3]:
$S_{\!EE}^{\rm univ}(R)$ $=$ $2\pi\ln\!\frac{R}{\epsilon}\!\int_{\Sigma_{\perp}}\!\sqrt{h}\,\langle T^{\mu}{}_{\mu}\rangle$ (1)
$=$  $\!\!\,2\pi\,\ln\!\frac{R}{\epsilon}\int_{r=R}\!R^{2}\,\mathrm{d}\Omega\;\langle T^{\mu}{}_{\mu}\rangle,$
as established by general arguments in relativistic QFT [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. [see Eq. (S13) of Appendix for a detailed derivation of Eq. (1).] In the infinite‑momentum frame (IMF), the two‑sphere $S^{2}_{R}$ projects onto the transverse (impact‑parameter) plane, and the surface measure reduces to the standard light‑front impact‑parameter measure used to define parton densities [15, 16].We use this mapping solely to express densities; the underlying entangling surface remains $S^{2}_{R}$.
By “mechanical entropy” we mean the universal logarithmic term $S_{\!EE}^{\rm univ}$ evaluated under two physical renormalization conditions: (i) a maximal‑entropy radius $R=R_{\!EE}$ set by the hadron; and (ii) a Lorentz‑contracted cutoff $\epsilon=R_{\!EE}e^{-y}$ for boost rapidity $y$. With these choices, spherical symmetry on $\Sigma_{\perp}$ gives $\int_{\Sigma_{\perp}}\!\sqrt{h}\,\langle T^{\mu}{}_{\mu}\rangle=4\pi R_{\!EE}^{2}\,\rho_{h}(R_{\!EE})$, so that
$S_{\!EE}(y)=c_{h}\,y,\qquad c_{h}=8\pi^{2}R_{\!EE}^{2}\rho_{h}(R_{\!EE}).$ (2)
At $y=0$ the universal logarithmic contribution is tuned to vanish by construction since $\epsilon=R_{\!EE}$ (therefore, our renormalization choice is valid only for large-$y$); this does not imply that the full state is pure.
For complementary 1+1D derivations of the linear‑in‑$y$ growth of entanglement, see [17, 18, 19, 20]. In particular, Ref. [20] computes the energy-momentum tensor (EMT) of a “hadron‑like” primary of dimension $h$ and finds (their Eq. 3.15) in pure 1+1D CFT $T_{++}(x_{+})=C_{h}/x_{+}^{2}$ with $C_{h}=h$; in the mass‑deformed case (their Eq. 3.19) one reads off the same coefficient from $C_{h}=x_{+}^{2}T_{++}(x_{+})$ at $x_{+}\!\to\!0,\infty$. This gives a universal slope $S_{EE}^{\rm univ}(y)=2\pi\,C_{h}\,y$, implying the dictionary $C_{h}=c_{h}/(2\pi)$ when compared to Eq. (2). Our result is the 3+1D QCD realization: we determine $c_{h}$ non‑perturbatively from the hadron’s trace anomaly localized on the entangling two‑sphere, $c_{h}=8\pi^{2}R_{EE}^{2}\rho_{h}(R_{EE})$, using lattice‑QCD gravitational form factor (GFF) inputs in 2D impact-parameter space Eq. (9).
It is convenient to parameterize the slope $c_{h}$ in terms of an operational (non‑perturbative) occupation number on the information wall:
$\frac{c_{h}}{2\pi}\ \stackrel{{\scriptstyle\text{op.}}}{{=}}\ \int_{0}^{1}\!dx\int d^{2}k_{\perp}\int d^{2}b_{\perp}\,\delta_{\Sigma}(b_{\perp})\,n_{q+g}(x,\bm{k}_{\perp},\bm{b}_{\perp})\,$, (3)
where $\delta_{\Sigma}(b_{\perp})\equiv\frac{\delta(b_{\perp}-R_{\!EE})}{2\pi R_{\!EE}}$, and $n_{q+g}(x,\bm{k_{\perp}},\bm{b}_{\perp})$ denotes the total (sea‑quark + gluon) occupation number with longitudinal momentum fraction $x$ and transverse momentum $\bm{k_{\perp}}$ per unit transverse phase‑space area $d^{2}b_{\perp}$; this identification serves as a non‑perturbative analogue of the gluon occupation number in the CGC/BFKL framework [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. Also note that $c_{h}$ per nucleon shows rapid saturation with increasing mass number $A$ in nuclei; see Fig. 1.
Formally, we define the hadron’s entanglement by starting from the vacuum density operator $|0\rangle\!\langle 0|$ and tracing out all quark‑ and gluon‑field modes whose support lies outside a sphere of radius $R$. The reduced density matrix admits the Euclidean path‑integral representation [2]
$\hat{\rho}_{<}(R)\;=\;\mathrm{Tr}_{r>R}\,|0\rangle\!\langle 0|={1\over\mathcal{N}}\int_{\begin{subarray}{c}\phi(\mathcal{C}_{>})=\phi_{>}\\ \phi(\mathcal{C}_{<})=\phi_{<}\end{subarray}}\mathcal{D}\phi\,e^{-I(\phi)}\,$, (4)
[…]
Percent-level lattice QCD calculations of the scalar (trace) gravitational form factors (GFFs) [39, 40, 41, 42] provide a direct, non-perturbative handle on $\langle T^{\mu}_{\ \mu}\rangle$. Independent experimental extractions corroborate these values: (i) the quark-sector GFF is obtained from deep-virtual Compton scattering at Jefferson Lab [43], analyzed with the dispersive–factorization approaches of [44, 45, 46]; (ii) the gluon-sector GFF is deduced from near-threshold $J/\psi$ photoproduction [47] using both holographic-QCD [48, 49] and an NRQCD–GPD factorization scheme [50].
The sign‑changing entropy gradient is an intrinsic property of the QCD vacuum, akin to Casimir-like stresses in non-Abelian gauge theories [51], and not a by‑product of the $2^{++}$$0^{++}$ resonances admixture that governs the Breit‑frame three‑dimensional EMT pressure inside a hadron [52, 53, 54, 55, 56, 57] (see Fig. S2 in the Appendix for the separate mechanical‑pressure node).
An entropic ansatz relates an elastic $2\!\to\!2$ cross section to the mechanical entropy of the color field produced in the collision [58, 59]; see Cross Sections and Data below for details. Ref. [60], following the exponential entropic ansatz of Ref. [19] for DIS structure functions, models total cross sections as $\sigma\propto e^{\delta y}$. Although this form reproduces vector–meson photoproduction data at high energies, it fails near threshold and overshoots the Froissart–Martin unitarity bound in the $y\!\to\!\infty$ limit. By contrast, our power‑law ansatz $\sigma\propto y^{\delta}$ describes both threshold and high‑energy photoproduction data with a single exponent and remains well below the Froissart–Martin limit across the full kinematic range, as illustrated in Figs. 4a4c. See also [61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73] for other proposals which relate entanglement entropy with multiplicity of final-state hadrons produced in high-energy collisions.
In the remainder of this Letter we (i) summarize the lattice‑QCD inputs that fix $\rho_{h}$ and $R_{\!EE}$; (ii) quantify the entropic force associated with $\partial_{R}S_{\!EE}$; and (iii) show how unitarity reduces an entropic ansatz for elastic $2\!\to\!2$ processes to $\sigma(s)=\mathcal{N}^{2}y^{\delta}$, with $\delta$ reflecting the UV sensitivity of the channel.
[…]
Nucleon ($J=\tfrac{1}{2}$).— With $t=(P^{\prime}-P)^{2}$, $\bar{P}=(P^{\prime}+P)/2$,
$\langle P^{\prime}|T_{i}^{\mu\nu}|P\rangle$ $=\bar{u}(P^{\prime})\!\Bigl[A_{i}(t)\gamma^{(\mu}\bar{P}^{\nu)}+D_{i}(t)\frac{q^{\mu}q^{\nu}-g^{\mu\nu}q^{2}}{4m_{N}}\Bigr]u(P),$ (5)
with $i\equiv\text{sea-quark}~(q)~\text{or gluon}~(g)$ and the trace is governed by the dipole
$A_{i}^{S}(t)=A_{i}(t)-\tfrac{3t}{4m_{N}^{2}}D_{i}(t)=\alpha_{S,i}(1-t/\Lambda_{S,i}^{2})^{-2}.$ (6)
Pion (J=0J=0).— With monopole form
$A_{\pi}^{S}(t)=A_{\pi}(t)-\tfrac{t}{4m_{\pi}^{2}}\bigl[2A_{\pi}(t)+3D_{\pi}(t)\bigr].$ (7)
Spin‑0 nuclei.— With $n_{A}=3A-1$,
$A_{i}^{S,A}(t)=A\,\alpha_{S,i}\,(1-t/\Lambda_{S,i}^{2})^{-n_{A}}.$ (8)
After 2D Fourier transform with respect to $t=-\bm{\Delta}_{\perp}^{2}$, any $n$-pole form yields the Bessel–$K$ density
$\rho_{i}(b_{\perp})=\alpha_{S,i}\frac{\Lambda_{S,i}^{2}}{2\pi}\frac{(\Lambda_{S,i}b_{\perp})^{n-1}}{2^{n-1}\Gamma(n)}K_{n-1}(\Lambda_{S,i}b_{\perp}).$ (9)
Summing over $i=q,g$ and evaluating at $b_{\perp}=R_{\!EE}$ fixes $c_{h}=8\pi^{2}R_{\!EE}^{2}\rho_{h}(R_{\!EE})$ in Eq. (2). All lattice fit parameters are given in Table S1 of Appendix.
Figure 1: Mechanical entanglement entropy per nucleon, $S_{\!EE}/(A\,y)$, versus entangling radius $R$. The light band spans $2\!\leq\!A\!\leq\!500$; dashed curves highlight representative nuclei. All curves level off at $S_{\!EE}/(A\,y)\!\simeq\!8.94$, signaling rapid saturation with $A$.
Figure 2: Radial derivative $\partial_{R}S_{\!EE}$ for the systems in Fig. 1. Positive values ($R<R_{\!EE}$) represent an outward entropic pressure; negative values ($R>R_{\!EE}$) pull inward. The zero crossing defines the “information wall” at $R_{\!EE}$.
Table 1: Charge radii $r_{c}$, maximum-entropy radii $R_{\!EE}$, and specific entropies $S_{\!EE}/(A\,y)$ for representative hadronic and nuclear systems. All radii are in fm. Proton $r_{c}$ is from PRad [81]; the pion value is the PDG average [82]; nuclear radii are taken from [83]. Over the mass range shown the maximum-entropy radius is reproduced to better than $0.1\,\%$ by $R_{\!EE}(A)=0.964\,\sqrt{A}\;-\;0.279/\sqrt{A}\;-\;0.014$. Note that $R_{\!EE}(A)$ need not match the liquid–drop $r_{0}A^{1/3}$ law for geometric radii because the underlying weights (2D trace anomaly density vs 3D mass density) and the codimension (surface vs volume) are different.
Species $r_{c}$ $R_{\!EE}$ $S_{\!EE}/(Ay)$
$\pi^{\pm}$ 0.659(4) 0.98(6) 12(4)
NN 0.831(14) 0.67(8) 7.17(36)
$^{4}\mathrm{He}$ 1.68(3) 1.77(21) 8.57(43)
$^{7}\mathrm{Li}$ 2.44(4) 2.43(28) 8.74(44)
$^{40}\mathrm{Ar}$ 3.43(3) 6.04(71) 8.91(45)
$^{208}\mathrm{Pb}$ 5.50(1) 13.9(1.6) 8.94(45)
Figure 3: Visual comparison of proton radii using lattice gravitational form factor (5) from [40], see Table S1: maximum-entropy radii $R_{\!EE}$, mechanical core $R_{q,g}$ (S18) (from the node of the 3D Breit‑frame pressure density, see Fig. S2), rms scalar radii $r_{s}$ (S20), rms energy radii $r_{00}$ (S20), and the experimental rms charge radius $r_{c}$ from PRad [81]. All rms radii shown are computed separately for the quark and for the gluon sectors using 3D Breit-frame densities and Eq. (S20) in the Appendix; no q+gq\!+\!g sum is taken. However, $R_{\!EE}$ is computed from $q+g$ 2D (impact-parameter space) trace anomaly density Eq. (9).
(a) $\phi$ photoproduction
(b) $J/\psi$ photoproduction
(c) $\Upsilon(1S)$ photoproduction
(d) Elastic pppp
Figure 4: Cross sections $\sigma$ are plotted versus the rapidity $y=\operatorname{arcosh}\gamma=\ln\!\bigl(\gamma+\sqrt{\gamma^{2}-1}\bigr)$, with $\gamma_{\text{beam}}\equiv E_{\gamma}/E_{\gamma,\mathrm{thr}}(V)$ for photoproduction and $\gamma_{\text{CM}}\equiv\sqrt{s}/(2m_{N})$ for elastic pppp scattering. Panels display (a) $\phi$ photoproduction, (b) $J/\psi$ photoproduction, (c) $\Upsilon(1S)$ photoproduction, and (d) elastic pppp scattering. The solid blue curves show the scaling ansatz $\sigma\propto y^{\delta}$ with $\delta_{\phi}=0.387$ in (a) and $\delta=2$ in (b)–(d). The shallower $y^{0.39}$ rise for $\phi$ indicates enhanced infrared sensitivity of light–quark dynamics, whereas heavy quarkonia and elastic pppp retain the full $y^{2}$ growth expected from the high-energy eikonal picture. Data sources: $\phi$ from Refs. [84, 85]; $J/\psi$ and $\Upsilon$ from the compilation in Ref. [48]; elastic pppp from Refs. [86, 87, 88].
Mechanical Entanglement and Confinement.— Equation (2), together with Figs. 1 and 2, shows that the entropy gradient reverses sign at $r=R_{\!EE}$ (see Table 1 for values of $R_{\!EE}$ for pion and nuclei, also see Fig. 3 for a visual comparison of various proton radii inferred from the lattice GFFs). Accordingly, the radial derivative $\partial_{R}S_{\!EE}$ acts as an entropic pressure or force in the sense of Refs. [89, 90]. A parton that strays inside the surface ($r<R_{\!EE}$) is pushed outward, whereas one that moves outside ($r>R_{\!EE}$) is pulled back. Colored degrees of freedom are therefore locked onto the two-dimensional “information wall” at $r=R_{\!EE}$. This entropic bag complements the Wilson area‑law criterion: it provides a high‑energy (UV) diagnostic while the area law governs long‑distance (IR) dynamics.
Cross Sections and Data.— The entropic ansatz ties an elastic $2\!\to\!2$ cross section to the entropy of the intermediate color field [59]. In the standard eikonal representation, unitarity gives
$\sigma_{\mathrm{el}}(s)=\!\int d^{2}b_{\perp}\,\bigl|1-e^{i\chi(b_{\perp},y)}\bigr|^{2},$ (10)
and, because the phase accumulates additively with rapidity, we take an entropy‑controlled eikonal $\chi(b_{\perp},y)\propto y$. This yields a single‑scale power law at high energy.
Consequently, for each elastic/exclusive channel, we write
$\sigma_{\cal C}(s)=\mathcal{N}_{\cal C}^{2}\,y^{\delta_{\cal C}},\qquad y=\operatorname{arcosh}\gamma,$ (11)
where $\mathcal{N}_{\cal C}$ is a single normalization constant per elastic/exclusive channel $\mathcal{C}$ and the exponent $\delta_{\cal C}$ quantifies how efficiently the process probes the ultraviolet anomaly density on the information wall [59].
UV–dominated channels.— Rapidity is defined as $y=\operatorname{arcosh}\gamma=\ln\bigl(\gamma+\sqrt{\gamma^{2}-1}\bigr)$. For vector–meson photoproduction with $s=W^{2}=\left(p_{\gamma}+p_{p}\right)^{2}$, we adopt the dimensionless beam–energy ratio $\gamma_{\text{beam}}\equiv E_{\gamma}/E_{\gamma,\mathrm{thr}}(V)$, where $E_{\gamma,\mathrm{thr}}(V)=\bigl[(m_{N}+M_{V})^{2}-m_{N}^{2}\bigr]/(2m_{N})$, so that $\gamma_{\text{beam}}=1$ at threshold and $y_{\text{beam}}=\operatorname{arcosh}\gamma_{\text{beam}}$. For elastic $pp$ scattering (with $\sqrt{s}=W$) in the c.m. frame the Lorentz factor of either proton is $\gamma_{\text{CM}}\equiv\sqrt{s}/(2m_{N})$, yielding $y_{\text{CM}}=\operatorname{arcosh}\gamma_{\text{CM}}$. Reactions characterized by compact transverse sizes (elastic $pp$, $J/\psi$, and $\Upsilon$ production) sample the full ultraviolet content of the wall; their exponents saturate at the unitarity limit $\delta=2$:
[…]
Even at the largest measured rapidities ($y_{\!\text{CM}}\!\simeq\!10$ at the LHC) the resulting cross sections remain below $32~\mathrm{mb}$, smaller by a factor $\gtrsim 2$ than the canonical $\pi/m_{\pi}^{2}\simeq 62.8~\mathrm{mb}$ scale; see Figs. 4b4d.
IR-sensitive channel.— The more diffuse $\phi$ meson samples the wall less efficiently, yielding a softer rise
$\delta_{\phi}=0.387,\qquad\mathcal{N}_{\phi}^{2}=3.3\times 10^{-4}~\text{mb},$
so that $\sigma_{\gamma p\to\phi p}$ stays far below $\sigma_{\!{}_{\mathrm{FM}}}$ across the entire energy range from threshold to multi-TeV photoproduction; see Fig. 4a.
All extracted cross sections remain well below the canonical Froissart–Martin scale $(\pi/m_{\pi}^{2})\ln^{2}\!s$ for $\sigma_{\mathrm{tot}}$ (with $\pi/m_{\pi}^{2}\simeq 62.8~\mathrm{mb}$ after unit conversion): elastic $pp$ lies by factors $\gtrsim 2$ below at present energies, whereas heavy‑quarkonia photoproduction is suppressed by $\sim 10\!-\!10^{3}$.
Conclusions.— We have shown that the QCD trace anomaly, quantified with percent-level lattice determinations of the scalar gravitational form factors, fixes a transverse “information wall” of radius $R_{\!EE}$. Tracing over quark and gluon modes outside that surface yields a finite, parameter-free mechanical entropy $S_{\!EE}=c_{h}\,y$ with $c_{h}$ fixed entirely by lattice inputs—whose sign-changing radial derivative locks colored fields to the wall at all momentum scales. This entropic mechanism provides a high‑energy diagnostic that complements the traditional Wilson‑loop area law: it remains consistent with asymptotic freedom in the ultraviolet and with linear confinement in the infrared. Imposing unitarity on an entropic ansatz for the amplitude yields $\sigma(s)\propto y^{\delta}$. Hard processes ($pp$, $J/\psi$, $\Upsilon$) saturate the unitarity limit $\delta=2$, exhibiting the familiar $y^{2}\!\sim\!\ln^{2}\!s$ growth pattern, while the more diffuse $\phi$ channel selects $\delta_{\phi}=0.387$. All extracted cross sections remain well below the canonical Froissart–Martin bound $(\pi/m_{\pi}^{2})\ln^{2}\!s$ for $\sigma_{\mathrm{tot}}$.
Outlook.— The entropic framework developed here delivers a quantitative baseline against which future high-precision data can be benchmarked. Upcoming measurements at the HL-LHC, Jefferson Lab (CLAS12, GlueX, and SoLID), and the Electron–Ion Collider will confront its predictions across an unprecedented kinematic range. Any systematic deviation—for example from odderon exchange, Regge exchanges, exotic multiquark resonances—will stand out cleanly against the fully quantified entropy background, offering a sharpened path to discovering genuinely new QCD dynamics.
Acknowledgements.
K.M. was supported by the U.S. Department of Energy under Grant No. DE-FG02-04ER41302. Additional support for this work was provided by the DOE Office of Science through Contract No. DE-AC05-06OR23177, under which Jefferson Science Associates, LLC operates the Thomas Jefferson National Accelerator Facility.
Appendix
[…]
Figure S1: Region AA is bounded by the entangling surface $\Sigma_{\perp}$ (the inner circle is a 1D illustration of the 2D surface); its complement is region $B$. The arrow denotes the radial coordinate $r$ from the center of AA to $\Sigma_{\perp}$ localized at $r=R=R_{\!EE}$.
Derivation of Eq. (1).— We work on the $t=0$ hypersurface and write $x^{\mu}=(t,r_{\parallel},\bm{r}_{\perp})$ with $r_{\parallel}\equiv x^{1}\in\mathbb{R}$ the longitudinal coordinate and $\bm{r}_{\perp}\equiv(x^{2},x^{3})\in\mathbb{R}^{2}$ the transverse vector ($\bm{x}\equiv(r_{\parallel},\bm{r}_{\perp})$). Its magnitude is $r_{\perp}=|\bm{r}_{\perp}|=b_{\perp}$ which can be identified with the impact-parameter space, and the three–dimensional radius is $r=\sqrt{r_{\parallel}^{2}+r_{\perp}^{2}}$. The volume element factorizes into longitudinal and transverse parts,
$\mathrm{d}^{3}r=\mathrm{d}r_{\parallel}\,\mathrm{d}^{2}r_{\perp},\qquad\mathrm{d}^{2}r_{\perp}=r_{\perp}\,\mathrm{d}r_{\perp}\,\mathrm{d}\phi,$
while in spherical coordinates $(r,\theta,\phi)$ $\mathrm{d}^{3}r=r^{2}\sin\theta\,\mathrm{d}r\,\mathrm{d}\theta\,\mathrm{d}\phi$.
The flat metric in longitudinal–transverse coordinates is
$\mathrm{d}s^{2}=-\,\mathrm{d}t^{2}+\mathrm{d}r_{\parallel}^{2}+\mathrm{d}r_{\perp}^{2}+r_{\perp}^{2}\,\mathrm{d}\phi^{2},$ (S1)
and, with the substitutions $r_{\perp}=r\sin\theta$ and $r_{\parallel}=r\cos\theta$, this becomes the familiar spherical‑coordinate form
$\mathrm{d}s^{2}=-\,\mathrm{d}t^{2}+\mathrm{d}r^{2}+r^{2}\bigl(\mathrm{d}\theta^{2}+\sin^{2}\theta\,\mathrm{d}\phi^{2}\bigr).$ (S2)
In a neighborhood of the entangling surface $\Sigma_{\perp}$, shown in Fig. S1, an arbitrary smooth metric admits the expansion
$ds^{2}=g_{\mu\nu}dx^{\mu}dx^{\nu}=-dt^{2}+dr^{2}+\bigl[h_{ab}(x_{\perp})+\mathcal{O}(r^{2})\bigr]dx_{\perp}^{a}dx_{\perp}^{b},$ (S3)
[…]
1. Modular Hamiltonian after tracing out the exterior. Tracing over all quark‑ and gluon‑field modes with support outside a sphere of radius $R$, as in Fig. S1, produces the reduced state $\hat{\rho}_{<}(R)=\mathrm{Tr}_{r>R}|0\rangle\!\langle 0|$. Its generator of modular flow, $K(R)\equiv-\ln\hat{\rho}_{<}(R)$, governs time evolution for an observer confined to $r<R$. For the planar entangling surface that bounds the right Rindler wedge $\mathcal{R}^{+}=\{x^{1}>|t|\}$[6], the Bisognano–Wichmann theorem [91, 92] identifies the modular Hamiltonian with the boost generator,
$K_{\mathcal{R}^{+}}=2\pi\!\int\!\mathrm{d}^{3}x\,\bigl[x^{1}T^{00}-tT^{01}\bigr].$
Setting $t=0$ one finds
$K_{\mathcal{R}^{+}}=2\pi\!\int_{x^{1}>0}\!\mathrm{d}r_{\parallel}\,\mathrm{d}^{2}r_{\perp}\;r_{\parallel}\,T_{00}(0;r_{\parallel},\bm{r}_{\perp}).$ (S4)
Because $K$ couples linearly to $T_{00}$, the eigenvalue density of $\hat{\rho}_{<}$ (the entanglement spectrum) is fixed by $\langle T_{00}\rangle$. In the Lorentz‑invariant vacuum with the mostly‑plus Minkowski metric, $T_{00}=-\tfrac{1}{4}T^{\mu}{}_{\mu}$; tracing Lorentz indices therefore ties the spectrum directly to the trace spectral density $\rho_{h}$.
2. Conformal map to a ball. A special conformal transformation with $a^{\mu}=(0,0,0,1/R)$ maps the half‑space $r_{\parallel}>0$ to the ball $r<R$ and rescales the spherical measure $r^{2}\sin\theta$:
$\bigl|\partial(r_{\parallel},\bm{r}_{\perp})/\partial(r,\theta,\phi)\bigr|=\Bigl(\tfrac{2R}{R^{2}+r^{2}}\Bigr)^{4}.$
Because $T_{00}$ has conformal weight 4, this Jacobian cancels and one obtains the sphere‑local modular Hamiltonian [13, 14]
$K(R)=2\pi\!\!\int_{r<R}\!\!\mathrm{d}r_{\parallel}\,\mathrm{d}^{2}r_{\perp}\;\frac{R^{2}-r^{2}}{2R}\;T_{00}(0;r_{\parallel},\bm{r}_{\perp}).$ (S5)
Evaluating its vacuum expectation value and tracing indices projects $T_{00}$ onto $T^{\mu}{}_{\mu}$, reproducing the spectral density $\rho_{h}(r_{\perp})$ that drives the mechanical entanglement entropy.
3. Contact term. Energy–momentum conservation, $\partial_{\mu}T^{\mu\nu}=0$, implies the Ward identity
$\partial_{\mu}\bigl\langle T^{\mu\nu}(x)\,T_{00}(y)\bigr\rangle=-\delta^{(3)}(\bm{x}-\bm{y})\,\partial_{y}^{\nu}\!\bigl\langle T_{00}(y)\bigr\rangle.$
Integrating the $\nu=0$ component over $t$ at equal times isolates a purely local distributional contribution. Tracing Lorentz indices on the first operator fixes the unique contact piece of the correlator [3],
$\langle T(0,\bm{x})\,T_{00}(0;r_{\parallel}^{\prime},\bm{r}_{\perp}^{\prime})\rangle$ $=\langle T(0,\bm{x})\rangle\,\delta(r_{\parallel}-r_{\parallel}^{\prime})\,\delta^{(2)}(\bm{r}_{\perp}-\bm{r}_{\perp}^{\prime})\,\,+\,\,\text{regular terms for }x\neq y.$ (S6)
4. Surface–local two‑point function. Inserting (S6) into (S5) gives
$\langle T(0,\bm{x}^{\!*})\,K\rangle$ $=2\pi\!\!\int_{r^{\prime 2}<R^{2}}\!\!\!\!\!\mathrm{d}r_{\parallel}^{\prime}\,\mathrm{d}^{2}r^{\prime}_{\perp}\;\frac{R^{2}-r_{\parallel}^{\prime 2}-r_{\perp}^{\prime 2}}{2R}\,T(0,\bm{x}^{\!*})\,\delta(r_{\parallel}^{\!*}\!-r_{\parallel}^{\prime})\,\delta^{(2)}(\bm{r}_{\perp}^{\!*}\!-\bm{r}_{\perp}^{\prime})$
$=2\pi\,T(0,\bm{x})\,F(r)\bigl|_{r=r^{\!*}},$ (S7)
where $F(r)=w(r)\,\Theta(R-r)$ with $w(r)=\tfrac{1}{2R}(R^{2}-r^{2})$.
For any spherically symmetric kernel $F(r)$,
$\nabla^{2}F(r)=\frac{1}{r^{2}}\frac{\partial}{\partial r}\!\Bigl(r^{2}\,\frac{\partial F}{\partial r}\Bigr),$
so $\tfrac{4\pi R^{2}}{2}\nabla^{2}$ projects the weight function onto the surface term dictated by Gauss’ theorem, isolating the local contact piece on the entangling sphere $r=R$. Differentiating yields
$\tfrac{4\pi R^{2}}{2}\nabla^{2}F(r)=\underbrace{\frac{3\times 4\pi R^{2}}{2R}\,\Theta(R-r)}_{\text{bulk term}}+\underbrace{\delta(R-r)}_{\delta_{\Sigma}(r)}.$ (S8)
The bulk contribution integrates to a power divergence (e.g. area‑law $\propto R^{2}/\epsilon^{2}$, curvature terms on $\Sigma_{\perp}$, etc.) removable by counter‑terms and therefore does not affect the universal logarithmic coefficient. The coefficient of $\delta(R-r)$ is unity by construction (Gauss’ theorem). Using (S8) in (S7) and keeping only the surface term,
$\langle T(0,\bm{x})\,K\rangle=2\pi\,T(0,\bm{x})\,\delta(R-r).$ (S9)
The factor $\delta(R-r)\equiv\delta_{\Sigma}(r)$ is a surface delta function normalized to $\int\!\mathrm{d}^{3}r\,\delta_{\Sigma}(r)=4\pi R^{2}$: it collapses any 3D integral onto the 2D entangling sphere $r=R$. Here $\delta_{\Sigma}^{(3D)}(r)=\delta(R-r)$ so $\int d^{3}r\,\delta_{\Sigma}^{(3D)}=4\pi R^{2}$, while in the main text (3) we use $\delta_{\Sigma}^{(2D)}(b_{\perp})=\delta(b_{\perp}-R)/(2\pi R)$ so $\int d^{2}b_{\perp}\,\delta_{\Sigma}^{(2D)}=1$. Concretely,
$\int\!\mathrm{d}^{3}r\,\delta_{\Sigma}(r)\,f(r,\Omega)=R^{2}\!\int_{r=R}\!\mathrm{d}\Omega\,f(R,\Omega)=\int_{\Sigma_{\perp}}\!\sqrt{h}\,f,$ (S10)
so only the values of the integrand on $\Sigma_{\perp}$ contribute.
Although Eq. (S9) was established within a conformal field theory, the derivation relies solely on locality and the Ward identity $\partial_{\mu}T^{\mu\nu}=0$; it therefore applies to any relativistic quantum field theory, including QCD. We thus adopt (S9) as the starting point for the explicit calculation of the entanglement–entropy flow that culminates in Eq. (S11).
5. Entanglement‑entropy flow. A uniform Weyl shift $\delta g_{\mu\nu}=2\sigma g_{\mu\nu}$ gives [2]
$\delta S_{\!EE}=-\sigma\!\int\!\mathrm{d}^{3}x\,\langle T(0,\bm{x})\,K\rangle=-\sigma\,2\pi\!\int_{r=R}\!\mathrm{d}^{2}r_{\perp}\;T(0;r_{\parallel},\bm{r}_{\perp}),$ (S11)
where we used (S9) to collapse the 3D integral to 2D. Choosing $\sigma=\delta R/R$ gives the RG‐flow equation
$R\,\frac{\mathrm{d}S_{\!EE}}{\mathrm{d}R}=-\,2\pi\!\int_{r=R}\!\mathrm{d}^{2}r_{\perp}\;T\bigl(0;r_{\parallel},\bm{r}_{\perp}\bigr).$ (S12)
Integrating from $\epsilon$ to $R$, defining $J(R^{\prime})=\int_{r=R^{\prime}}\mathrm{d}^{2}r_{\perp}\,T(0;r_{\parallel},\bm{r}_{\perp})$, and discarding the analytic (counter‑term) piece yields the universal logarithm:
$\boxed{S_{\!EE}^{\text{univ}}(R)=\,2\pi\,\ln\!\frac{R}{\epsilon}\int_{r=R}\!\mathrm{d}^{2}r_{\perp}\;T\bigl(0;r_{\parallel},\bm{r}_{\perp}\bigr)=\,2\pi\,\ln\!\frac{R}{\epsilon}\int_{r=R}\!R^{2}\,\mathrm{d}\Omega\;T\bigl(0;R,\Omega\bigr)=\,2\pi\,\ln\!\frac{R}{\epsilon}\int_{\Sigma_{\perp}}R^{2}\,\!\mathrm{d}\theta\mathrm{d}\phi\sqrt{h}\,T\bigl(0;R,\Omega\bigr),\,}$ (S13)
[…]
$\left.\mathrm{d}s^{2}\right|_{t=0,r=R}=R^{2}h_{ab}dx^{a}dx^{b}=R^{2}\bigl(\mathrm{d}\theta^{2}+\sin^{2}\theta\,\mathrm{d}\phi^{2}\bigr)\,.$ (S14)
with its determinant $h=\sin^{2}\theta$.
Mechanical core radius and scalar/energy rms radii from lattice GFF.— We work with the sea-quark and gluon contributions to the proton EMT, $T_{q,g}^{\mu\nu}$, whose matrix elements are parameterized by the gravitational form factors (GFFs) as in Eq. (5), with $t=(P^{\prime}-P)^{2}=-K^{2}$. Using the scalar–tensor separation of GFFs, the scalar ($0^{++}$) combination is
$A^{S}_{q,g}(t)\equiv A_{q,g}(t)-\frac{3t}{4m_{N}^{2}}\,D_{q,g}(t)\,$, (S15)
while $A_{q,g}(t)$ carries the traceless ($2^{++}$) part. Throughout this section, all radii quoted and plotted in Fig. 3 of the main text are extracted separately for quarks and for gluons; no q+gq\!+\!g sum is taken when computing the individual radii.
Mechanical core radius from the Breit-frame pressure node
The isotropic pressure density in the Breit frame is
$p_{q,g}(r)=\frac{1}{6m_{N}\,r^{2}}\frac{d}{dr}\!\left[r^{2}\frac{d}{dr}D_{q,g}(r)\right],$ (S16)
and obeys the stability (von Laue) condition $\int_{0}^{\infty}\!dr\,4\pi r^{2}\,p_{q,g}(r)=0$. The mechanical core radius $R_{q,g}$ is defined by the node of the surface force $4\pi r^{2}p_{q,g}(r)$, i.e. $p_{q,g}(R_{q,g})=0$, where the repulsive ($2^{++}$) and attractive ($0^{++}$) stresses balance, see Fig. S2.
For the two–scale dipole representation used in Table S1,
$A_{q,g}(t)=\frac{\alpha_{A,i}}{(1-t/\Lambda_{A,q/g}^{2})^{2}},\qquad A^{S}_{q,g}(t)=\frac{\alpha_{S,i}}{(1-t/\Lambda_{S,q/g}^{2})^{2}},$
(with $\alpha_{S,i}\!\simeq\!\alpha_{A,i}$ in the lattice fits), the corresponding pressure admits a simple analytic form, [^1] This closed form follows from the standard 3D Fourier transform of a dipole, together with $D_{q,g}(t)=\tfrac{4m_{N}^{2}}{3t}\!\left[A_{q,g}(t)-A^{S}_{q,g}(t)\right]$.
$p_{q,g}(r)=\frac{\alpha_{A,i}\,m_{N}}{36\pi}\Bigl[\Lambda_{A,q/g}^{3}\,e^{-\Lambda_{A,q/g}r}-\Lambda_{S,q/g}^{3}\,e^{-\Lambda_{S,q/g}r}\Bigr],$ (S17)
which satisfies the von Laue condition identically. The node condition $p_{q,g}(R_{q,g})=0$ gives the mechanical core radius in closed form,
$R_{q,g}=\frac{3\,\ln(\Lambda_{A,q/g}/\Lambda_{S,q/g})}{\Lambda_{A,q/g}-\Lambda_{S,q/g}}\;,\quad\text{(in }\mathrm{GeV}^{-1}\text{; multiply by }\hbar c=0.19733~\text{fm/GeV)}.$ (S18)
The values used in Fig. 3 are obtained by inserting the dipole masses $(\Lambda_{A,q/g},\Lambda_{S,q/g})$ from Table S1.
Figure S2: Radial distribution $r^{2}p_{q,g}(r)$ normalized to $A_{q,g}(0)$, computed from Eq. (S17). The node marks the mechanical equilibrium radius $R_{q,g}$ where repulsive ($2^{++}$) and attractive ($0^{++}$) stresses balance.
Scalar and energy rms radii (quark/gluon separately)
Definitions.
The slope of the Lorentz–invariant scalar form factor $A^{S}_{q,g}(t)\equiv\eta_{\mu\nu}T^{\mu\nu}_{q,g}(t)$ at the origin defines a frame–independent rms radius, while the slope of the energy density $A^{00}_{q,g}(t)=T^{00}_{q,g}(t)$—taken in the nucleon rest frame—defines a distinct Breit–frame radius:
$\bigl\langle r^{2}_{s,q(g)}\bigr\rangle$ $=-6\,\left.\frac{dA^{S}_{q,g}(t)}{dK^{2}}\right|_{K^{2}=0},\quad\text{with }t=-K^{2},$
$\bigl\langle r^{2}_{00,q(g)}\bigr\rangle$ $=-6\,\left.\frac{dA^{00}_{q,g}(t)}{dK^{2}}\right|_{K^{2}=0},$ (S19)
where
$A^{S}_{q,g}(t)$ $=A_{q,g}(t)-\frac{3t}{4m_{N}^{2}}\,D_{q,g}(t),$
$A^{00}_{q,g}(t)$ $=A_{q,g}(t)-\frac{t}{4m_{N}^{2}}\,D_{q,g}(t).$
Two–scale dipole result.
Within the two–scale dipole model used in Table S1,   $A^{S}_{q,g}(t)=\alpha_{S,i}/(1-t/\Lambda_{S,q/g}^{2})^{2}$, $A_{q,g}(t)=\alpha_{A,i}/(1-t/\Lambda_{A,q/g}^{2})^{2}$, and $D_{q,g}(t)=\tfrac{4m_{N}^{2}}{3t}\!\left[A_{q,g}(t)-A^{S}_{q,g}(t)\right]$. For $\alpha_{S,i}=\alpha_{A,i}$ (as in current lattice fits) one finds the compact analytic expressions
$\boxed{\ \langle r^{2}_{s,q(g)}\rangle=\frac{12}{\Lambda_{S,q/g}^{2}},\qquad\langle r^{2}_{00,q(g)}\rangle=\frac{8}{\Lambda_{A,q/g}^{2}}+\frac{4}{\Lambda_{S,q/g}^{2}}\,}$ (S20)
in $\mathrm{GeV}^{-2}$; multiply by $\hbar^{2}c^{2}=(0.19733~\mathrm{fm})^{2}$ to convert to $\mathrm{fm}^{2}$. These formulas are applied separately to the quark and to the gluon sectors using $\Lambda_{A,q/g},\Lambda_{S,q/g}$ from Table S1, yielding the individual rms radii displayed in Fig. 3.
Lattice GFF fit parameters.—Here we summarize the lattice fit parameters for the gravitational form factors (GFF) used in the main text.
Table S1: Dipole $A_{X}(t)$] and $D$-term $D_{X}(t)$] fit parameters from lattice-QCD analyses [40, 41]. The first block lists nucleon quark ($q$) and gluon ($g$) sectors; the second block shows the corresponding pion sectors. Masses are in GeV and uncertainties are statistical. Scalar form factors are constructed as $A_{N}^{S}=A_{q}^{S}+A_{g}^{S}$ for the nucleon and $A_{\pi}^{S}=A_{q}^{S}+A_{g}^{S}$ at $m_{\pi}=170~\text{MeV}$. For completeness, the nucleon scalar-pole masses (that reproduce the lattice D-term) are $\Lambda_{S,q}=0.804(108)\,\text{GeV}$ and $\Lambda_{S,g}=0.596(77)\,\text{GeV}$.
Species Sector $\alpha_{A}=\alpha_{S}$ $\Lambda_{A}$ $\alpha_{D}$ $\Lambda_{D}$
Nucleon $q$ 0.510(25) 1.477(44) 1.30(49)-1.30(49) 0.81(14)
$g$ 0.501(27) 1.262(18) 2.57(84)-2.57(84) 0.538(65)
Pion $q$ 0.481(15) 1.262(37) 0.304(26)-0.304(26) 1.44(21)
$g$ 0.546(18) 1.129(41) 0.596(65)-0.596(65) 0.677(65)

### v4

@@
 Abstract
-We formulate confinement in QCD as an entropic surface phenomenon.
-Quark and gluon quantum information is localized on a transverse, entangling two‑sphere of radius $R_{\!EE}$; at this radius the QCD vacuum—partitioned by a hadron into interior and exterior regions—reaches its maximal entanglement entropy.
-Lattice‑QCD determinations of the scalar (trace) gravitational form factors fix both $R_{\!EE}$ and the transverse trace‑anomaly density $\rho_{h}(R_{\!EE})$, yielding a parameter‑free slope $c_{h}=8\pi^{2}R_{\!EE}^{2}\rho_{h}(R_{\!EE})$ and a mechanical entropy $S_{\!EE}(y)=c_{h}\,y$ that grows linearly with rapidity $y$.
-The entropy gradient $\partial_{R}S_{\!EE}$ changes sign at $R_{\!EE}$: it pushes colored degrees of freedom outward for $r<R_{\!EE}$ and pulls them inward for $r>R_{\!EE}$, thereby localizing them on the codimension‑2 entangling two‑sphere $\Sigma_{\perp}=S^{2}_{R_{\!EE}}$ (which, in the infinite‑momentum frame (IMF), projects onto the transverse plane)—the “information wall.” This provides a high‑energy (large‑$y$) entropic confinement diagnostic that complements—rather than replaces—Wilson’s area‑law criterion, which probes long‑distance dynamics near the rest frame ($y\!\to\!0$).
+We formulate confinement in quantum chromodynamics (QCD) as an entropic surface phenomenon.
+Quark and gluon quantum information is localized on a transverse, entangling two-sphere of radius $R_{\!EE}$; at this radius the QCD vacuum—partitioned by a hadron into interior and exterior regions—reaches its maximal entanglement entropy.
+Lattice-QCD determinations of the scalar (trace) gravitational form factors fix both $R_{\!EE}$ and the transverse trace-anomaly density $\rho_{h}(R_{\!EE})$, yielding a parameter-free slope $c_{h}=8\pi^{2}R_{\!EE}^{2}\rho_{h}(R_{\!EE})$ and a mechanical entropy $S_{\!EE}(y)=c_{h}\,y$ that grows linearly with rapidity $y$.
+The entropy gradient $\partial_{R}S_{\!EE}$ changes sign at $R_{\!EE}$: it pushes colored degrees of freedom outward for $r<R_{\!EE}$ and pulls them inward for $r>R_{\!EE}$, thereby localizing them on the codimension-2 entangling two-sphere $\Sigma_{\perp}=S^{2}_{R_{\!EE}}$ (which, in the infinite-momentum frame (IMF), projects onto the transverse plane)—the “information wall.” This provides a high-energy (large-$y$) entropic confinement diagnostic that complements—rather than replaces—Wilson’s area-law criterion, which probes long-distance dynamics near the rest frame ($y\!\to\!0$).
 Imposing unitarity on an entropic ansatz for the amplitude yields $\sigma(s)\!\propto\!y^{\delta}$.
-World data favor $\delta=2$ for elastic $pp\,(p\bar{p})$ scattering and heavy‑quark photoproduction, whereas $\phi$ photoproduction favors a softer $\delta=0.387$.
+World data favor $\delta=2$ for elastic $pp\,(p\bar{p})$ scattering and heavy-quark photoproduction, whereas $\phi$ photoproduction favors a softer $\delta=0.387$.
 All extracted cross sections remain well below the Froissart–Martin bound.
-These results provide a confinement criterion quantified directly from non‑perturbative QCD inputs, unifying the trace anomaly, entanglement entropy, and high‑energy scattering within a single quantitative framework.
+These results provide a confinement criterion quantified directly from nonperturbative QCD inputs, unifying the trace anomaly, entanglement entropy, and high-energy scattering within a single quantitative framework.
 Introduction.— Why do colored quarks and gluons never emerge from hadrons?
-The textbook answer invokes an area law for large Wilson loops, which diagnoses a linear confining potential in the IR [1].
+The textbook answer invokes an area law for large Wilson loops, which diagnoses a linear confining potential in the infrared (IR) [1].
 Yet that criterion carries no information about ultraviolet (UV) dynamics, where asymptotic freedom reigns.
 Here we propose a complementary, UV–sensitive diagnostic: entropic confinement.
-Quantum information carried by quarks and gluons localizes on a codimension‑2 entangling surface of fixed radius $R_{\!EE}$.
+Quantum information carried by quarks and gluons localizes on a codimension-2 entangling surface of fixed radius $R_{\!EE}$.
 Geometrically, the entangling surface is the two–sphere $\Sigma_{\perp}=S^{2}_{R}$ equipped with the induced metric $h_{ab}$ and area element $\sqrt{h}\,d^{2}x$.
-The universal contribution to the von Neumann entropy of the reduced state is then controlled by the QCD trace anomaly on $\Sigma_{\perp}$ [2, 3]:
+The universal contribution to the von Neumann entropy of the reduced state is then controlled by the quantum chromodynamics (QCD) trace anomaly on $\Sigma_{\perp}$ [2, 3]:
 $S_{\!EE}^{\rm univ}(R)$ $=$ $2\pi\ln\!\frac{R}{\epsilon}\!\int_{\Sigma_{\perp}}\!\sqrt{h}\,\langle T^{\mu}{}_{\mu}\rangle$ (1)
@@
 as established by general arguments in relativistic QFT [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14].
-[see Eq.
-(S13) of Appendix for a detailed derivation of Eq.
-(1).] In the infinite‑momentum frame (IMF), the two‑sphere $S^{2}_{R}$ projects onto the transverse (impact‑parameter) plane, and the surface measure reduces to the standard light‑front impact‑parameter measure used to define parton densities [15, 16].We use this mapping solely to express densities; the underlying entangling surface remains $S^{2}_{R}$.
-By “mechanical entropy” we mean the universal logarithmic term $S_{\!EE}^{\rm univ}$ evaluated under two physical renormalization conditions: (i) a maximal‑entropy radius $R=R_{\!EE}$ set by the hadron; and (ii) a Lorentz‑contracted cutoff $\epsilon=R_{\!EE}e^{-y}$ for boost rapidity $y$.
+See the Supplemental Material (included at the end of this arXiv version) for a detailed derivation of Eq.
+(1).
+In the infinite-momentum frame (IMF), the two-sphere $S^{2}_{R}$ projects onto the transverse (impact-parameter) plane, and the surface measure reduces to the standard light-front impact-parameter measure used to define parton densities [15, 16].
+We use this mapping solely to express densities; the underlying entangling surface remains $S^{2}_{R}$.
+By “mechanical entropy” we mean the universal logarithmic term $S_{\!EE}^{\rm univ}$ evaluated under two physical renormalization conditions: (i) a maximal-entropy radius $R=R_{\!EE}$ set by the hadron; and (ii) a Lorentz-contracted cutoff $\epsilon=R_{\!EE}e^{-y}$ for boost rapidity $y$.
 With these choices, spherical symmetry on $\Sigma_{\perp}$ gives $\int_{\Sigma_{\perp}}\!\sqrt{h}\,\langle T^{\mu}{}_{\mu}\rangle=4\pi R_{\!EE}^{2}\,\rho_{h}(R_{\!EE})$, so that
@@
 At $y=0$ the universal logarithmic contribution is tuned to vanish by construction since $\epsilon=R_{\!EE}$ (therefore, our renormalization choice is valid only for large-$y$); this does not imply that the full state is pure.
-For complementary 1+1D derivations of the linear‑in‑$y$ growth of entanglement, see [17, 18, 19, 20].
+For complementary 1+1D derivations of the linear-in-$y$ growth of entanglement, see [17, 18, 19, 20].
 In particular, Ref.
-[20] computes the energy-momentum tensor (EMT) of a “hadron‑like” primary of dimension $h$ and finds (their Eq.
-3.15) in pure 1+1D CFT $T_{++}(x_{+})=C_{h}/x_{+}^{2}$ with $C_{h}=h$; in the mass‑deformed case (their Eq.
+[20] computes the energy-momentum tensor (EMT) of a “hadron-like” primary of dimension $h$ and finds (their Eq.
+3.15) in pure 1+1D CFT $T_{++}(x_{+})=C_{h}/x_{+}^{2}$ with $C_{h}=h$; in the mass-deformed case (their Eq.
 3.19) one reads off the same coefficient from $C_{h}=x_{+}^{2}T_{++}(x_{+})$ at $x_{+}\!\to\!0,\infty$.
@@
 (2).
-Our result is the 3+1D QCD realization: we determine $c_{h}$ non‑perturbatively from the hadron’s trace anomaly localized on the entangling two‑sphere, $c_{h}=8\pi^{2}R_{EE}^{2}\rho_{h}(R_{EE})$, using lattice‑QCD gravitational form factor (GFF) inputs in 2D impact-parameter space Eq.
+Our result is the 3+1D QCD realization: we determine $c_{h}$ nonperturbatively from the hadron’s trace anomaly localized on the entangling two-sphere, $c_{h}=8\pi^{2}R_{EE}^{2}\rho_{h}(R_{EE})$, using lattice-QCD gravitational form factor (GFF) inputs in 2D impact-parameter space Eq.
 (9).
-It is convenient to parameterize the slope $c_{h}$ in terms of an operational (non‑perturbative) occupation number on the information wall:
-$\frac{c_{h}}{2\pi}\ \stackrel{{\scriptstyle\text{op.}}}{{=}}\ \int_{0}^{1}\!dx\int d^{2}k_{\perp}\int d^{2}b_{\perp}\,\delta_{\Sigma}(b_{\perp})\,n_{q+g}(x,\bm{k}_{\perp},\bm{b}_{\perp})\,$, (3)
-where $\delta_{\Sigma}(b_{\perp})\equiv\frac{\delta(b_{\perp}-R_{\!EE})}{2\pi R_{\!EE}}$, and $n_{q+g}(x,\bm{k_{\perp}},\bm{b}_{\perp})$ denotes the total (sea‑quark + gluon) occupation number with longitudinal momentum fraction $x$ and transverse momentum $\bm{k_{\perp}}$ per unit transverse phase‑space area $d^{2}b_{\perp}$; this identification serves as a non‑perturbative analogue of the gluon occupation number in the CGC/BFKL framework [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38].
+It is convenient to parametrize the slope $c_{h}$ in terms of an operational (nonperturbative) occupation number on the information wall:
+$\frac{c_{h}}{8\pi^{2}R_{\!EE}^{2}\,}\ \stackrel{{\scriptstyle\text{op.}}}{{=}}\ \int_{0}^{1}\!dx\int d^{2}k_{\perp}\int d^{2}b_{\perp}\,\delta_{\Sigma}(b_{\perp})\,n_{q+g}(x,\bm{k}_{\perp},\bm{b}_{\perp})\,$, (3)
+where $\delta_{\Sigma}(b_{\perp})\equiv\frac{\delta(b_{\perp}-R_{\!EE})}{2\pi R_{\!EE}}$, and $n_{q+g}(x,\bm{k_{\perp}},\bm{b}_{\perp})$ denotes the total (sea-quark + gluon) occupation number with longitudinal momentum fraction $x$ and transverse momentum $\bm{k_{\perp}}$ per unit transverse phase-space area $d^{2}b_{\perp}$; this identification serves as a nonperturbative analog of the gluon occupation number in the CGC/BFKL framework [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38].
 Also note that $c_{h}$ per nucleon shows rapid saturation with increasing mass number $A$ in nuclei; see Fig.
 1.
-Formally, we define the hadron’s entanglement by starting from the vacuum density operator $|0\rangle\!\langle 0|$ and tracing out all quark‑ and gluon‑field modes whose support lies outside a sphere of radius $R$.
-The reduced density matrix admits the Euclidean path‑integral representation [2]
+Formally, we define the hadron’s entanglement by starting from the vacuum density operator $|0\rangle\!\langle 0|$ and tracing out all quark- and gluon-field modes whose support lies outside a sphere of radius $R$, see Fig.
+S1 in the Supplemental Material.
+The reduced density matrix admits the Euclidean path-integral representation [2]
 $\hat{\rho}_{<}(R)\;=\;\mathrm{Tr}_{r>R}\,|0\rangle\!\langle 0|={1\over\mathcal{N}}\int_{\begin{subarray}{c}\phi(\mathcal{C}_{>})=\phi_{>}\\ \phi(\mathcal{C}_{<})=\phi_{<}\end{subarray}}\mathcal{D}\phi\,e^{-I(\phi)}\,$, (4)
 […]
-Percent-level lattice QCD calculations of the scalar (trace) gravitational form factors (GFFs) [39, 40, 41, 42] provide a direct, non-perturbative handle on $\langle T^{\mu}_{\ \mu}\rangle$.
+Percent-level lattice QCD calculations of the scalar (trace) gravitational form factors (GFFs) [39, 40, 41, 42] provide a direct, nonperturbative handle on $\langle T^{\mu}_{\ \mu}\rangle$.
 Independent experimental extractions corroborate these values: (i) the quark-sector GFF is obtained from deep-virtual Compton scattering at Jefferson Lab [43], analyzed with the dispersive–factorization approaches of [44, 45, 46]; (ii) the gluon-sector GFF is deduced from near-threshold $J/\psi$ photoproduction [47] using both holographic-QCD [48, 49] and an NRQCD–GPD factorization scheme [50].
-The sign‑changing entropy gradient is an intrinsic property of the QCD vacuum, akin to Casimir-like stresses in non-Abelian gauge theories [51], and not a by‑product of the $2^{++}$$0^{++}$ resonances admixture that governs the Breit‑frame three‑dimensional EMT pressure inside a hadron [52, 53, 54, 55, 56, 57] (see Fig.
-S2 in the Appendix for the separate mechanical‑pressure node).
+The sign-changing entropy gradient (see Fig.
+2) is an intrinsic property of the QCD vacuum, akin to Casimir-like stresses in non-Abelian gauge theories [51], and not a by-product of the $2^{++}$$0^{++}$ resonances admixture that governs the Breit-frame three-dimensional EMT pressure inside a hadron [52, 53, 54, 55, 56, 57] (see Fig.
+S2 in the Supplemental Material for the separate mechanical-pressure node).
 An entropic ansatz relates an elastic $2\!\to\!2$ cross section to the mechanical entropy of the color field produced in the collision [58, 59]; see Cross sections and data below for details.
@@
 [19] for DIS structure functions, models total cross sections as $\sigma\propto e^{\delta y}$.
-Although this form reproduces vector–meson photoproduction data at high energies, it fails near threshold and overshoots the Froissart–Martin unitarity bound in the $y\!\to\!\infty$ limit.
-By contrast, our power‑law ansatz $\sigma\propto y^{\delta}$ describes both threshold and high‑energy photoproduction data with a single exponent and remains well below the Froissart–Martin limit across the full kinematic range, as illustrated in Figs.
+Although this form reproduces vector-meson photoproduction data at high energies, it fails near threshold and overshoots the Froissart–Martin unitarity bound in the $y\!\to\!\infty$ limit.
+By contrast, our power-law ansatz $\sigma\propto y^{\delta}$ describes both threshold and high-energy photoproduction data with a single exponent and remains well below the Froissart–Martin limit across the full kinematic range, as illustrated in Figs.
 4a4c.
 See also [61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73] for other proposals which relate entanglement entropy with multiplicity of final-state hadrons produced in high-energy collisions.
-In the remainder of this Letter we (i) summarize the lattice‑QCD inputs that fix $\rho_{h}$ and $R_{\!EE}$; (ii) quantify the entropic force associated with $\partial_{R}S_{\!EE}$; and (iii) show how unitarity reduces an entropic ansatz for elastic $2\!\to\!2$ processes to $\sigma(s)=\mathcal{N}^{2}y^{\delta}$, with $\delta$ reflecting the UV sensitivity of the channel.
+In the remainder of this paper we (i) summarize the lattice-QCD inputs that fix $\rho_{h}$ and $R_{\!EE}$; (ii) quantify the entropic force associated with $\partial_{R}S_{\!EE}$; and (iii) show how unitarity reduces an entropic ansatz for elastic $2\!\to\!2$ processes to $\sigma(s)=\mathcal{N}^{2}y^{\delta}$, with $\delta$ reflecting the UV sensitivity of the channel.
 […]
@@
 $A_{\pi}^{S}(t)=A_{\pi}(t)-\tfrac{t}{4m_{\pi}^{2}}\bigl[2A_{\pi}(t)+3D_{\pi}(t)\bigr].$ (7)
-Spin‑0 nuclei.— With $n_{A}=3A-1$,
+Spin-0 nuclei.— With $n_{A}=3A-1$,
 $A_{i}^{S,A}(t)=A\,\alpha_{S,i}\,(1-t/\Lambda_{S,i}^{2})^{-n_{A}}.$ (8)
@@
 $\rho_{i}(b_{\perp})=\alpha_{S,i}\frac{\Lambda_{S,i}^{2}}{2\pi}\frac{(\Lambda_{S,i}b_{\perp})^{n-1}}{2^{n-1}\Gamma(n)}K_{n-1}(\Lambda_{S,i}b_{\perp}).$ (9)
-Summing over $i=q,g$ and evaluating at $b_{\perp}=R_{\!EE}$ fixes $c_{h}=8\pi^{2}R_{\!EE}^{2}\rho_{h}(R_{\!EE})$ in Eq. (2). All lattice fit parameters are given in Table S1 of Appendix.
+Summing over $i=q,g$ and evaluating at $b_{\perp}=R_{\!EE}$ fixes $c_{h}=8\pi^{2}R_{\!EE}^{2}\rho_{h}(R_{\!EE})$ in Eq.
+(2).
+All lattice fit parameters are summarized in Table S1 of the Supplemental Material.
 Figure 1: Mechanical entanglement entropy per nucleon, $S_{\!EE}/(A\,y)$, versus entangling radius $R$.
@@
 $^{208}\mathrm{Pb}$ 5.50(1) 13.9(1.6) 8.94(45)
-Figure 3: Visual comparison of proton radii using lattice gravitational form factor (5) from [40], see Table S1: maximum-entropy radii $R_{\!EE}$, mechanical core $R_{q,g}$ (S18) (from the node of the 3D Breit‑frame pressure density, see Fig.
+Figure 3: Visual comparison of proton radii using lattice gravitational form factor (5) from [40], see Table S1: maximum-entropy radius $R_{\!EE}$, mechanical core $R_{q,g}$ (S18) (from the node of the 3D Breit-frame pressure density, see Fig.
 S2), rms scalar radii $r_{s}$ (S20), rms energy radii $r_{00}$ (S20), and the experimental rms charge radius $r_{c}$ from PRad [81].
 All rms radii shown are computed separately for the quark and for the gluon sectors using 3D Breit-frame densities and Eq.
-(S20) in the Appendix; no q+gq\!+\!g sum is taken.
+(S20) in the Supplemental Material; no q+gq\!+\!g sum is taken.
 However, $R_{\!EE}$ is computed from $q+g$ 2D (impact-parameter space) trace anomaly density Eq.
@@
 Colored degrees of freedom are therefore locked onto the two-dimensional “information wall” at $r=R_{\!EE}$.
-This entropic bag complements the Wilson area‑law criterion: it provides a high‑energy (UV) diagnostic while the area law governs long‑distance (IR) dynamics.
+This entropic bag complements the Wilson area-law criterion: it provides a high-energy (UV) diagnostic while the area law governs long-distance (IR) dynamics.
 Cross sections and data.— The entropic ansatz ties an elastic $2\!\to\!2$ cross section to the entropy of the intermediate color field [59]. In the standard eikonal representation, unitarity gives
 $\sigma_{\mathrm{el}}(s)=\!\int d^{2}b_{\perp}\,\bigl|1-e^{i\chi(b_{\perp},y)}\bigr|^{2},$ (10)
-and, because the phase accumulates additively with rapidity, we take an entropy‑controlled eikonal $\chi(b_{\perp},y)\propto y$. This yields a single‑scale power law at high energy.
-Consequently, for each elastic/exclusive channel, we write
+and, because the phase accumulates additively with rapidity, we take an entropy-controlled eikonal $\chi(b_{\perp},y)\propto y$.
+This yields a single-scale power law at high energy.
+In particular, for small-eikonal (“moderate opacity”) regime, we find $\sigma_{\mathrm{el}}(s)\propto\int d^{2}b\,|\chi|^{2}\propto y^{2}$.
+Consequently, for each elastic/exclusive channel, we write the entropic ansatz
 $\sigma_{\cal C}(s)=\mathcal{N}_{\cal C}^{2}\,y^{\delta_{\cal C}},\qquad y=\operatorname{arcosh}\gamma,$ (11)
@@
 UV–dominated channels.— Rapidity is defined as $y=\operatorname{arcosh}\gamma=\ln\bigl(\gamma+\sqrt{\gamma^{2}-1}\bigr)$.
-For vector–meson photoproduction with $s=W^{2}=\left(p_{\gamma}+p_{p}\right)^{2}$, we adopt the dimensionless beam–energy ratio $\gamma_{\text{beam}}\equiv E_{\gamma}/E_{\gamma,\mathrm{thr}}(V)$, where $E_{\gamma,\mathrm{thr}}(V)=\bigl[(m_{N}+M_{V})^{2}-m_{N}^{2}\bigr]/(2m_{N})$, so that $\gamma_{\text{beam}}=1$ at threshold and $y_{\text{beam}}=\operatorname{arcosh}\gamma_{\text{beam}}$.
+For vector-meson photoproduction with $s=W^{2}=\left(p_{\gamma}+p_{p}\right)^{2}$, we adopt the dimensionless beam–energy ratio $\gamma_{\text{beam}}\equiv E_{\gamma}/E_{\gamma,\mathrm{thr}}(V)$, where $E_{\gamma,\mathrm{thr}}(V)=\bigl[(m_{N}+M_{V})^{2}-m_{N}^{2}\bigr]/(2m_{N})$, so that $\gamma_{\text{beam}}=1$ at threshold and $y_{\text{beam}}=\operatorname{arcosh}\gamma_{\text{beam}}$.
 For elastic $pp$ scattering (with $\sqrt{s}=W$) in the c.m.
@@
 so that $\sigma_{\gamma p\to\phi p}$ stays far below $\sigma_{\!{}_{\mathrm{FM}}}$ across the entire energy range from threshold to multi-TeV photoproduction; see Fig. 4a.
-All extracted cross sections remain well below the canonical Froissart–Martin scale $(\pi/m_{\pi}^{2})\ln^{2}\!s$ for $\sigma_{\mathrm{tot}}$ (with $\pi/m_{\pi}^{2}\simeq 62.8~\mathrm{mb}$ after unit conversion): elastic $pp$ lies by factors $\gtrsim 2$ below at present energies, whereas heavy‑quarkonia photoproduction is suppressed by $\sim 10\!-\!10^{3}$.
+All extracted cross sections remain well below the canonical Froissart–Martin scale $(\pi/m_{\pi}^{2})\ln^{2}\!s$ for $\sigma_{\mathrm{tot}}$ (with $\pi/m_{\pi}^{2}\simeq 62.8~\mathrm{mb}$ after unit conversion): elastic $pp$ lies by factors $\gtrsim 2$ below at present energies, whereas heavy-quarkonia photoproduction is suppressed by $\sim 10\!-\!10^{3}$.
 Conclusions.— We have shown that the QCD trace anomaly, quantified with percent-level lattice determinations of the scalar gravitational form factors, fixes a transverse “information wall” of radius $R_{\!EE}$.
 Tracing over quark and gluon modes outside that surface yields a finite, parameter-free mechanical entropy $S_{\!EE}=c_{h}\,y$ with $c_{h}$ fixed entirely by lattice inputs—whose sign-changing radial derivative locks colored fields to the wall at all momentum scales.
-This entropic mechanism provides a high‑energy diagnostic that complements the traditional Wilson‑loop area law: it remains consistent with asymptotic freedom in the ultraviolet and with linear confinement in the infrared.
+This entropic mechanism provides a high-energy diagnostic that complements the traditional Wilson-loop area law: it remains consistent with asymptotic freedom in the ultraviolet and with linear confinement in the infrared.
 Imposing unitarity on an entropic ansatz for the amplitude yields $\sigma(s)\propto y^{\delta}$.
@@
 Outlook.— The entropic framework developed here delivers a quantitative baseline against which future high-precision data can be benchmarked.
-Upcoming measurements at the HL-LHC, Jefferson Lab (CLAS12, GlueX, and SoLID), and the Electron–Ion Collider will confront its predictions across an unprecedented kinematic range.
+Upcoming measurements at the HL-LHC, Jefferson Lab (CLAS12, GlueX, and SoLID), and the Electron-Ion Collider will confront its predictions across an unprecedented kinematic range.
 Any systematic deviation—for example from odderon exchange, Regge exchanges, exotic multiquark resonances—will stand out cleanly against the fully quantified entropy background, offering a sharpened path to discovering genuinely new QCD dynamics.
@@
 DE-AC05-06OR23177, under which Jefferson Science Associates, LLC operates the Thomas Jefferson National Accelerator Facility.
-Appendix
+DATA AVAILABILITY
+The data that support the findings of this article are not publicly available. The data are available from the authors upon reasonable request.
+Supplemental Material
 […]
@@
 (1).— We work on the $t=0$ hypersurface and write $x^{\mu}=(t,r_{\parallel},\bm{r}_{\perp})$ with $r_{\parallel}\equiv x^{1}\in\mathbb{R}$ the longitudinal coordinate and $\bm{r}_{\perp}\equiv(x^{2},x^{3})\in\mathbb{R}^{2}$ the transverse vector ($\bm{x}\equiv(r_{\parallel},\bm{r}_{\perp})$).
-Its magnitude is $r_{\perp}=|\bm{r}_{\perp}|=b_{\perp}$ which can be identified with the impact-parameter space, and the three–dimensional radius is $r=\sqrt{r_{\parallel}^{2}+r_{\perp}^{2}}$.
+Its magnitude is $r_{\perp}=|\bm{r}_{\perp}|=b_{\perp}$ which can be identified with the impact-parameter space, and the three-dimensional radius is $r=\sqrt{r_{\parallel}^{2}+r_{\perp}^{2}}$.
 The volume element factorizes into longitudinal and transverse parts,
@@
 $\mathrm{d}s^{2}=-\,\mathrm{d}t^{2}+\mathrm{d}r_{\parallel}^{2}+\mathrm{d}r_{\perp}^{2}+r_{\perp}^{2}\,\mathrm{d}\phi^{2},$ (S1)
-and, with the substitutions $r_{\perp}=r\sin\theta$ and $r_{\parallel}=r\cos\theta$, this becomes the familiar spherical‑coordinate form
+and, with the substitutions $r_{\perp}=r\sin\theta$ and $r_{\parallel}=r\cos\theta$, this becomes the familiar spherical-coordinate form
 $\mathrm{d}s^{2}=-\,\mathrm{d}t^{2}+\mathrm{d}r^{2}+r^{2}\bigl(\mathrm{d}\theta^{2}+\sin^{2}\theta\,\mathrm{d}\phi^{2}\bigr).$ (S2)
@@
 Modular Hamiltonian after tracing out the exterior.
-Tracing over all quark‑ and gluon‑field modes with support outside a sphere of radius $R$, as in Fig.
+Tracing over all quark- and gluon-field modes with support outside a sphere of radius $R$, as in Fig.
 S1, produces the reduced state $\hat{\rho}_{<}(R)=\mathrm{Tr}_{r>R}|0\rangle\!\langle 0|$.
@@
 Because $K$ couples linearly to $T_{00}$, the eigenvalue density of $\hat{\rho}_{<}$ (the entanglement spectrum) is fixed by $\langle T_{00}\rangle$.
-In the Lorentz‑invariant vacuum with the mostly‑plus Minkowski metric, $T_{00}=-\tfrac{1}{4}T^{\mu}{}_{\mu}$; tracing Lorentz indices therefore ties the spectrum directly to the trace spectral density $\rho_{h}$.
-2. Conformal map to a ball. A special conformal transformation with $a^{\mu}=(0,0,0,1/R)$ maps the half‑space $r_{\parallel}>0$ to the ball $r<R$ and rescales the spherical measure $r^{2}\sin\theta$:
+In the Lorentz-invariant vacuum with the mostly-plus Minkowski metric, $T_{00}=-\tfrac{1}{4}T^{\mu}{}_{\mu}$; tracing Lorentz indices therefore ties the spectrum directly to the trace spectral density $\rho_{h}$.
+2. Conformal map to a ball. A special conformal transformation with $a^{\mu}=(0,0,0,1/R)$ maps the half-space $r_{\parallel}>0$ to the ball $r<R$ and rescales the spherical measure $r^{2}\sin\theta$:
 $\bigl|\partial(r_{\parallel},\bm{r}_{\perp})/\partial(r,\theta,\phi)\bigr|=\Bigl(\tfrac{2R}{R^{2}+r^{2}}\Bigr)^{4}.$
-Because $T_{00}$ has conformal weight 4, this Jacobian cancels and one obtains the sphere‑local modular Hamiltonian [13, 14]
+Because $T_{00}$ has conformal weight 4, this Jacobian cancels and one obtains the sphere-local modular Hamiltonian [13, 14]
 $K(R)=2\pi\!\!\int_{r<R}\!\!\mathrm{d}r_{\parallel}\,\mathrm{d}^{2}r_{\perp}\;\frac{R^{2}-r^{2}}{2R}\;T_{00}(0;r_{\parallel},\bm{r}_{\perp}).$ (S5)
@@
 $\langle T(0,\bm{x})\,T_{00}(0;r_{\parallel}^{\prime},\bm{r}_{\perp}^{\prime})\rangle$ $=\langle T(0,\bm{x})\rangle\,\delta(r_{\parallel}-r_{\parallel}^{\prime})\,\delta^{(2)}(\bm{r}_{\perp}-\bm{r}_{\perp}^{\prime})\,\,+\,\,\text{regular terms for }x\neq y.$ (S6)
-4. Surface–local two‑point function. Inserting (S6) into (S5) gives
+4. Surface-local two-point function. Inserting (S6) into (S5) gives
 $\langle T(0,\bm{x}^{\!*})\,K\rangle$ $=2\pi\!\!\int_{r^{\prime 2}<R^{2}}\!\!\!\!\!\mathrm{d}r_{\parallel}^{\prime}\,\mathrm{d}^{2}r^{\prime}_{\perp}\;\frac{R^{2}-r_{\parallel}^{\prime 2}-r_{\perp}^{\prime 2}}{2R}\,T(0,\bm{x}^{\!*})\,\delta(r_{\parallel}^{\!*}\!-r_{\parallel}^{\prime})\,\delta^{(2)}(\bm{r}_{\perp}^{\!*}\!-\bm{r}_{\perp}^{\prime})$
@@
 The bulk contribution integrates to a power divergence (e.g.
-area‑law $\propto R^{2}/\epsilon^{2}$, curvature terms on $\Sigma_{\perp}$, etc.) removable by counter‑terms and therefore does not affect the universal logarithmic coefficient.
+area-law $\propto R^{2}/\epsilon^{2}$, curvature terms on $\Sigma_{\perp}$, etc.) removable by counter-terms and therefore does not affect the universal logarithmic coefficient.
 The coefficient of $\delta(R-r)$ is unity by construction (Gauss’ theorem).
@@
 (S11).
-5. Entanglement‑entropy flow. A uniform Weyl shift $\delta g_{\mu\nu}=2\sigma g_{\mu\nu}$ gives [2]
+5. Entanglement-entropy flow. A uniform Weyl shift $\delta g_{\mu\nu}=2\sigma g_{\mu\nu}$ gives [2]
 $\delta S_{\!EE}=-\sigma\!\int\!\mathrm{d}^{3}x\,\langle T(0,\bm{x})\,K\rangle=-\sigma\,2\pi\!\int_{r=R}\!\mathrm{d}^{2}r_{\perp}\;T(0;r_{\parallel},\bm{r}_{\perp}),$ (S11)
-where we used (S9) to collapse the 3D integral to 2D. Choosing $\sigma=\delta R/R$ gives the RG‐flow equation
+where we used (S9) to collapse the 3D integral to 2D. Choosing $\sigma=\delta R/R$ gives the RG-flow equation
 $R\,\frac{\mathrm{d}S_{\!EE}}{\mathrm{d}R}=-\,2\pi\!\int_{r=R}\!\mathrm{d}^{2}r_{\perp}\;T\bigl(0;r_{\parallel},\bm{r}_{\perp}\bigr).$ (S12)
-Integrating from $\epsilon$ to $R$, defining $J(R^{\prime})=\int_{r=R^{\prime}}\mathrm{d}^{2}r_{\perp}\,T(0;r_{\parallel},\bm{r}_{\perp})$, and discarding the analytic (counter‑term) piece yields the universal logarithm:
+Integrating from $\epsilon$ to $R$, and discarding the analytic (counter-term) piece yields the universal logarithm:
 $\boxed{S_{\!EE}^{\text{univ}}(R)=\,2\pi\,\ln\!\frac{R}{\epsilon}\int_{r=R}\!\mathrm{d}^{2}r_{\perp}\;T\bigl(0;r_{\parallel},\bm{r}_{\perp}\bigr)=\,2\pi\,\ln\!\frac{R}{\epsilon}\int_{r=R}\!R^{2}\,\mathrm{d}\Omega\;T\bigl(0;R,\Omega\bigr)=\,2\pi\,\ln\!\frac{R}{\epsilon}\int_{\Sigma_{\perp}}R^{2}\,\!\mathrm{d}\theta\mathrm{d}\phi\sqrt{h}\,T\bigl(0;R,\Omega\bigr),\,}$ (S13)
@@
 Definitions.
-The slope of the Lorentz–invariant scalar form factor $A^{S}_{q,g}(t)\equiv\eta_{\mu\nu}T^{\mu\nu}_{q,g}(t)$ at the origin defines a frame–independent rms radius, while the slope of the energy density $A^{00}_{q,g}(t)=T^{00}_{q,g}(t)$—taken in the nucleon rest frame—defines a distinct Breit–frame radius:
+The slope of the Lorentz-invariant scalar form factor $A^{S}_{q,g}(t)\equiv\eta_{\mu\nu}T^{\mu\nu}_{q,g}(t)$ at the origin defines a frame-independent rms radius, while the slope of the energy density $A^{00}_{q,g}(t)=T^{00}_{q,g}(t)$—taken in the nucleon rest frame—defines a distinct Breit-frame radius:
 $\bigl\langle r^{2}_{s,q(g)}\bigr\rangle$ $=-6\,\left.\frac{dA^{S}_{q,g}(t)}{dK^{2}}\right|_{K^{2}=0},\quad\text{with }t=-K^{2},$